### v1

# Resolving the Bulk-Boundary Correspondence Paradox on Low-Symmetry Surfaces of Weyl Semimetals

Abstract
Low-symmetry surfaces of topological semimetals offer access to boundary phenomena hidden on conventional facets, yet systematic studies remain scarce due to experimental challenges and the lack of a general framework for reconciling bulk and surface periodicities[1, 2, 3, 4, 5]. Here, we investigate the (103) surface of the Weyl semimetal NdAlSi using angle resolved photoemission spectroscopy and density functional theory. The (103) surface is an example of a low symmetry surface that presents an apparent paradox to the bulk-boundary correspondence. The surface periodicity to which the topological surface states are expected to adhere does not correspond to the bulk periodicity of the Weyl points. By showing that successive bulk Brillouin zones generate replicas that accumulate into a superlattice commensurate with the surface Brillouin zone, we demonstrate how the apparent bulk boundary correspondence paradox is resolved and establish a universal criterion for arbitrary facets. The framework and experimental results further suggests that overlapping Fermi arcs can hybridize into closed Fermi arc loops, enriching boundary topology and enabling unconventional transport, interference, and collective phenomena unique to low-symmetry facets.
Introduction
Topological semimetals have emerged as a fertile ground for exploring exotic quantum states of matter, where nontrivial band topology gives rise to robust surface excitations and novel bulk transport phenomena[1, 2, 3]. Among them, Weyl semimetals are characterized by linear bulk band crossings–Weyl points–that act as monopoles of Berry curvature in momentum space[1, 2, 3, 6, 7, 8, 9, 10, 11]. A defining feature of Weyl semimetals is the bulk-boundary correspondence (BBC), which dictates that pairs of Weyl points of opposite chirality in the bulk are connected by open Fermi arc states on the surface[4, 5]. This correspondence has been extensively established on high-symmetry surfaces such as (001) and (100), where the surface Brillouin zone (BZ) is commensurate with the bulk projection[7, 8, 9, 12, 13, 14, 16, 15].
In contrast, low-symmetry surfaces remain largely unexplored, even though they are uniquely positioned to reveal fundamentally new physics. Their intrinsic incommensurability between surface and bulk BZ periodicities raises a conceptual puzzle: while Weyl points project according to the bulk BZ, surface Fermi arcs follow the surface BZ (SBZ), this mismatch suggesting an apparent breakdown of the BBC. At the same time, low-symmetry facets naturally disentangle surface states and surface-bulk projected states (SBPSs) from real bulk states. On high-symmetry surfaces, rotational and mirror symmetries enforce strong constraints and cancellations that suppress many boundary responses. By contrast, on low-symmetry surfaces these constraints are lifted: Fermi arcs can become highly anisotropic or undergo reconstruction[14, 17, 18], parallel arc segments can enhance nesting-driven instabilities[19, 20, 21, 22], Berry-curvature contributions no longer cancel and instead give rise to nonlinear and nonreciprocal responses[23, 24, 25, 26], and the reduction of crystalline symmetries facilitates the emergence of higher-order boundary states[27]. Despite this promise, experimental and theoretical studies of low-symmetry surfaces have been scarce, owing to the experimental difficulty of preparing and interrogating non-natural cleavage planes and the lack of a general theoretical framework for reconciling bulk and surface periodicities on such facets[1, 2, 3, 4, 5].
Here we address this challenge by investigating the low-symmetry (103) surface of the Weyl semimetal NdAlSi using angle resolved photoemission spectroscopy (ARPES) and density functional theory (DFT). By combining disorder-induced suppression of coherent surface states on low-symmetry (103) facets, we introduce a new experimental strategy that enables the direct and unambiguous identification of surface-bulk resonance states (SBRSs)[28, 29, 30, 31, 32, 33, 34, 35] in a Weyl semimetal. At the same time, we demonstrate that the apparent paradox of BBC is resolved when bulk-to-surface projections are extended beyond the first BZ: higher-order projections generate translated replicas whose superposition yields a superlattice periodicity commensurate with the SBZ. This establishes a generalized framework for BBC on arbitrary crystal facets, demonstrating that the apparent breakdown on low-symmetry surfaces is not a failure of the BBC but a consequence of incomplete bulk projection. The resulting least common multiple (LCM) criterion provides a universal means of reconciling bulk and surface states and opens a pathway for engineering exotic boundary phenomena across a broad class of topological materials. More broadly, our results open a new path in the study of low-symmetry surfaces of topological semimetals, laying the groundwork for systematic exploration of their hidden boundary physics. By showing how such facets naturally disentangle and amplify the interplay between surface and bulk states, our findings create opportunities to design unconventional responses and collective excitations that remain inaccessible on conventional high-symmetry surfaces.
Results
To address above issues, we turn to the Weyl semimetal NdAlSi[13, 14, 15, 22], a member of the non-centrosymmetric RAlSi (R=rare earth) family[36, 37, 38, 39] that provides an excellent platform for studying low-symmetry facets. In particular, we investigate the (103) surface, which is never explored yet offers a unique opportunity to directly test the validity of BBC. The crystallographic structure of NdAlSi and the orientation of the (103) surface are shown in Fig. 1a. This cut exposes a low-symmetry plane that is rarely studied in ARPES due to the experimental challenges associated with surface preparation. Fig. 1b illustrates the construction of the bulk BZ and its projection onto the (103) surface. Unlike high-symmetry surfaces such as (001) or (100), the (103) surface projection produces a fundamental mismatch: the periodicity of the SBZ (green lines) does not coincide with that of the projected bulk BZ (pink lines), as shown in Fig. 1c.
The surface orientation was verified by X-ray diffraction (XRD), which confirms the high-quality exposure of the (103) plane (Fig. 1d) (see Methods for sample preparation). Fig. 1e and 1f display Fermi surfaces of NdAlSi measured at 150 eV on well defined (Fig. 1e) and disordered (Fig. 1f) regions of the (103) surface. On the well defined area, the measured Fermi surface (Fig. 1e) is dominated by surface states (see Section 1 in Supplementary Materials for details). On the disordered region, however, the surface states are almost suppressed[15], revealing additional features that cannot be assigned to pure surface bands. Fig. 1e clearly shows that the surface states respect the periodicity of the SBZ. In contrast, Fig. 1f reveals that the bulk states exhibit a periodicity that does not coincide with the SBZ. This provides direct experimental evidence of the intrinsic non-commensurability between the surface states and bulk state projections on the (103) surface. While not unexpected for topologically trivial systems, this observation raises a central question for Weyl semimetals: If Weyl points follow the bulk BZ periodicity while surface Fermi arcs follow the SBZ periodicity, does the BBC break down on low-symmetry surfaces?
To understand the nature of the electronic states on the (103) surface and answer the above question, we performed systematic ARPES measurements combined with surface-projected DFT calculations based on Green’s function methods. Fig. 2a shows the Fermi surface measured at 41 eV photon energy on a well defined region of the surface, which is dominated by surface states (see Section 2 in Supplementary Materials for results under different polarizations). The corresponding DFT calculation for a Si-terminated surface (Fig. 2b) reproduces the observed features quite well, confirming their surface origin. Data collected on a disordered region of the surface under the same conditions show that the surface states are strongly suppressed (Fig. 2c)[15]. This suppression acts as a natural “filter”, allowing otherwise hidden SBPSs to emerge. When a surface state enters the bulk-projected region, it hybridizes with the bulk continuum and evolves into a SBRS[28, 29, 30, 31, 32, 33, 34, 35], resulting in hybridized features observable in ARPES (see Section 3 in the Supplementary Materials). The DFT calculation reproduces these features, as shown in Fig. 2b. Consistently, Green’s function calculations capture both sharp surface states and additional bulk-projected features.
Momentum-resolved band dispersions along two representative cuts (Cut1 and Cut2) are shown in Figs. 2e and 2f. These sharp dispersive bands agree well with surface-projected DFT calculations (Figs. 2g and 2h). By contrast, along the same cuts on the disordered surface (Figs. 2i and 2j), broader and bulk-like bands appear, consistent with calculated SBPSs (Figs. 2k and 2l). Sections 3 and 4 of the Supplementary Materials provides more details on the identification of surface states, bulk states, SBPSs and SBRSs.
Having established the surface states and SBPSs, we next turn to the bulk electronic structure underlying the (103) surface. The 3D bulk Fermi surface from DFT is shown in Fig. 3a, while Fig. 3b illustrates the (103) surface orientation and its perpendicular cross-section in momentum space. Calculated bulk Fermi surface slices along the [103] direction (k⟂) and perpendicular to the (103) surface (cyan plane in Fig. 3b) are presented in Figs. 3c and 3d, serving as theoretical benchmarks.
Experimentally, the photon energy dependent ARPES spectral intensity map at the Fermi level along ky (Fig. 3e) reveals strong k⟂ modulation, a hallmark of bulk-derived states. Systematic Fermi surface measurements over photon energies from 125 eV to 195 eV (Figs. 3f-i, 3n-q) further confirm this k⟂ dispersion. The corresponding DFT calculations (Figs. 3j-m, 3r-u) reproduce the experimental features with remarkable accuracy. More detailed bulk Fermi surfaces of the NdAlSi (103) surface are provided in Section 5 of the Supplementary Materials.
Beyond establishing the bulk band structure, these results provide critical insight into the apparent breakdown of the BBC. In particular, the evolution of bulk Fermi surface slices along [103] reveals a systematic in-plane translation of features as k⟂ is varied. Each slice produces a shifted replica of the bulk Fermi surface within the SBZ, reflecting the intrinsic periodicity of bulk states along the [103] direction. As successive bulk BZs are considered, these replicas accumulate into an effective superlattice periodicity that is distinct from the primitive bulk BZ but naturally commensurate with the SBZ.
From this perspective, the apparent mismatch between the periodicity of projected Weyl points and that of the surface Fermi arcs is a direct consequence of restricting the projection to the first bulk BZ. The bulk projections along [103] demonstrate experimentally that multiple BZs must be included: only then does the accumulated superlattice restore the periodicity needed to match the surface states. This insight provides the experimental foundation for resolving the BBC paradox, as developed below.
To clarify how this resolves the apparent BCC paradox, we track how the bulk-projection periodicity evolves as higher-order BZs are included. If only the first bulk BZ is considered, the projected Weyl points (Fig. 4a) clearly mismatch the theoretical SBZ (purple lines) and the experimental SBZ (green lines). Adding the second-order (Fig. 4b) and third-order BZs and Weyl points (Fig. 4c) along progressively restores the periodicity, bringing the bulk projection closer to the surface states. This restoration follows a LCM criterion: the projected displacement between successive bulk BZs closes only after an integer multiple of the in-plane reciprocal vectors. For the (103) surface of NdAlSi, the minimum closure occurs after steps (see Section 6 in the Supplementary Materials). As illustrated in Fig. 4d, the projected Weyl points recover the theoretical SBZ periodicity (purple lines) once higher-order BZs are included.
Beyond this restoration, the higher‐order projections reveal an emergent smaller repeating unit: a “new SBZ” (green lines in Fig. 4d) whose period along is one third of the theoretical SBZ. This “new SBZ” coincides with the one observed experimentally. Its origin lies in the fact that the bulk BZs of three consecutive, nonequivalent (103) surface unit cells differ by a phase of 2/3 in momentum space; their hybridization produces the “new smaller SBZ”. Equivalently, this effect can be viewed as a fixed lateral shift between successive bulk projections (see Section 6 in the Supplementary Materials), which manifests as a momentum-space moir pattern–a weak superlattice modulation with a period one third of the original SBZ. In general, whether a single bulk BZ suffices depends on surface symmetry. For high-symmetry surfaces such as (001), the first bulk BZ already coincides with the SBZ (Fig. 4f). For low-symmetry surfaces such as (103), however, successive bulk projections are laterally shifted (Fig. 4g), and only after steps does the bulk periodicity close. Exact spectral reconstruction is achieved only in the infinite-projection limit (see Section 6 in the Supplementary Materials). Thus, the observed mismatch is not a breakdown of BBC but rather the natural consequence of incomplete bulk projection, while the emergent new SBZ reflects the moir-like interference among multiple shifted projections.
This insight also clarifies why different computational schemes perform so differently. Finite-slab calculations (six layers in our case) truncate the bulk projections and therefore miss the SBPSs that are essential on low-symmetry facets. By contrast, semi-infinite surface Green’s function calculations include contributions from infinitely many bulk BZs and thereby capture both surface states and SBPSs in excellent agreement with experiment. Thus, while slab models may suffice for commensurate high-symmetry surfaces[13, 14], Green’s function methods are indispensable for correctly describing the electronic structure of low-symmetry surfaces such as (103) (see Section 4 in the Supplementary Materials).
An additional subtlety arises when considering the surface Fermi arcs themselves. On high-symmetry surfaces such as (001), arcs from successive bulk BZs project in perfect registry, perhaps typically preserving the canonical picture of open arcs that directly connect projected Weyl points (Fig. 4f). In contrast, on low-symmetry surfaces like (103), the incommensurate periodicities of bulk and SBZs cause consecutive arcs to project with mismatched positions. This misalignment naturally drives hybridization between arcs originating from different bulk zones (Fig. 4g). Consequently, the observed arcs are not simple replicas but collective superpositions, and in certain cases multiple arcs can merge into closed surface Fermi arc loops (SFALs)–a boundary topology inaccessible on high-symmetry facets (Fig. 4d, see Section 7 of the Supplementary Materials for details). While our framework and present ARPES superposition features suggest the emergence of SFALs, their unambiguous demonstration remains open. Future high-resolution spin ARPES or quantum oscillation studies will be key to confirming their presence. Such loop-like structures enrich the boundary phenomenology of Weyl semimetals: their closed topology provides natural -space trajectories for novel oscillatory and interference effects, while their overlapping geometry may promote nesting-driven instabilities or correlated orders.
By investigating the low-symmetry (103) surface of the Weyl semimetal NdAlSi, we introduce a new experimental route to directly identify SBRSs. This is enabled by disorder-induced suppression of surface states, which acts as a natural filter to distinguish the SBRS from others. Furthermore, the (103) surface highlights an apparent paradox of the BBC. Contrary to the topological expectation that Fermi arcs and Weyl points should share the same periodicity, the Weyl points follow the bulk while the Fermi arcs display the surface periodicity. We resolve this puzzle by showing that the mismatch originates not from a break down of the BBC but from incomplete bulk projection. Successive BZ Weyl points generate translated surface Fermi arcs whose superposition yields a superlattice commensurate with the surface, thereby restoring the BBC. This LCM construction can be viewed as a momentum-space moir modulation, the k-space analogue of real-space moir stripes, where long-period modulations govern the reconstruction of boundary states.
More broadly, our results inaugurate a new frontier in the study of low-symmetry surfaces. The possibility of SFALs enriches the boundary topology of Weyl semimetals, suggesting unconventional transport, interference, and collective phenomena beyond the canonical open-arc paradigm. The conceptual framework extends beyond Weyl semimetals to Dirac and nodal-line semimetals, higher-order topological systems, and even non-centrosymmetric superconductors. Its analogy to moir physics in stacked two-dimensional materials further highlights its generality: in both real space and momentum space, mismatched periodicities generate emergent superlattices that control hybridization and boundary states. The same principle also bridges to quasicrystals, where intrinsic incommensurability requires infinite projection to restore quasi-periodicity. A more detailed discussion is provided in Section 8 of the Supplementary Materials. In summary, our work resolves the apparent BBC paradox on low-symmetry surfaces and provides a unifying momentum-space moir framework connecting crystalline and quasicrystalline boundaries, opening a platform to uncover hidden boundary phenomena and engineer exotic responses in quantum matter.
Methods
Sample growth Single crystals of NdAlSi were grown from Al as flux. Nd, Al, Si elements were sealed in an alumina crucible with the molar ratio of 1 : 10 : 1. The crucible was finally sealed in a highly evacuated quartz tube. The tube was heated up to 1273 K, maintained for 12 hours and then cooled down to 973 K at a rate of 3 K per hour. Single crystals were separated from the flux by centrifuging. The Al flux attached to the single crystals were removed by dilute NaOH solution.
Sample preparation of the (103) surface To obtain the low-symmetry (103) surface of NdAlSi, we first determined the crystallographic orientation of the single crystals by single crystal XRD. Based on the identified orientation, the crystals were cut with a high-precision diamond wire saw to expose the desired (103) direction. Prior to mounting, a shallow pre-cut was introduced along the target orientation, which facilitated controlled cleavage of the sample inside the ARPES chamber. This procedure enabled reliable preparation of well-defined (103) surfaces suitable for spectroscopic measurements.
ARPES measurements High-resolution ARPES measurements were performed at the I05 beamline of the Diamond synchrotron and at the Bloch beamline of the MAX IV synchrotron. The total energy resolution (analyzer and beamline) was set at 1520 meV for the measurements. The angular resolution of the analyser was 0.1 degree. The beamline spot size on the sample was about 50 m50 m at the I05 beamline of the Diamond synchrotron and about 10 m12 m at the Bloch beamline of the MAX IV synchrotron. The samples were cleaved in situ and measured in ultrahigh vacuum with a base pressure better than 1.010-10 mbar at about 8 K at the I05 beamline of the Diamond synchrotron and about 18 K at the Bloch beamline of the MAX IV synchrotron.
DFT calculations The electronic structure calculations for NdAlSi were performed based on the density functional theory (DFT)[40, 41] as implemented in the VASP package[42, 43]. The exchange-correlation functional was treated within the generalized gradient approximation (GGA) in the form of Perdew–Burke–Ernzerhof (PBE)[44]. The projector augmented wave (PAW) method[45, 46] was used to describe the electron–ion interactions. Since the measurement temperature is above the Curie temperature, NdAlSi is non-magnetic; therefore, the Nd pseudopotential without 4f electrons was employed in the calculations. A plane-wave kinetic energy cutoff of 360 eV was used. Brillouin zone sampling was performed using a 6×6×6 Monkhorst–Pack k-point grid [47]. A Gaussian smearing width of 0.1 eV was applied to approximate the Fermi–Dirac distribution. Spin–orbit coupling (SOC) was included as a second-order perturbation.
For the (001) surface states of NdAlSi, a slab model with 36 atomic layers and a 20 Å vacuum layer was constructed. The calculated results are consistent with experimental data. For the (103) surface states and Fermi arcs, we employed an iterative Green’s function method to model a semi-infinite surface and compute its surface states. Among the four possible terminations of the (103) surface, the Si-Al plane termination yielded results consistent with experimental observations.
The bulk Fermi surface of NdAlSi was computed using a 60×60×60 k-point grid including SOC. The projection of bulk states onto the plane for different -space slices was calculated via the Green’s function method, using a Fermi broadening of 0.01 eV to obtain the spectral function.
Data Availability
The authors declare that all data supporting the findings of this study are available within the paper and its Supplementary Information files.
Acknowledgement
The work presented here was financially supported by the Knut and Alice Wallenberg foundation (2018.0104). Y.G.S. acknowledges the National Natural Science Foundation of China (Grants No. U2032204), the Informatization Plan of Chinese Academy of Sciences (CAS-WX2021SF-0102) and the National Key R&D Program of China 2024YFA1408400. Q. S. W. acknowledges the National Natural Science Foundation of China No. 12274436. We acknowledge MAX IV Laboratory for time on Beamline BLOCH under Proposal 20241395 and 20250284 as well as Diamond Light Source for time on Beamline I05 under Proposal SI39652. Research conducted at MAX IV, a Swedish national user facility, is supported by the Swedish Research council under contract 2018-07152, the Swedish Governmental Agency for Innovation Systems under contract 2018-04969, and Formas under contract 2019-02496.
Author Contributions
C.L. proposed and conceived the project. C.L. carried out the ARPES experiments with the assistance from M.H.B. and F.S.. Z.L.Y. contributed to the band structure calculations and theoretical discussions. H.X.L. and Y.G.S. contributed to NdAlSi crystal growth. C.L. and D.P. contributed to XRD measurements. C.L. developed the data analysis software, performed the data analysis, contributed to the theoretical derivation and wrote the paper. T.K., J.O. and B.T. provided the beamline support. C.L., Z.L.Y., M.H.B., D.P., J.F.Z., T.X., Q.S.W. and O.T. participate in the scientific discussions. O.T. revised the manuscript. All authors participated in and commented on the paper.
Competing Interests
The authors declare no competing interests.