### v3

# Hidden Moiré Topology of Low-Symmetry Weyl Surfaces

Abstract
Topological materials are defined by the correspondence between bulk topology and boundary states, yet this correspondence becomes enigmatic on low-symmetry surfaces where bulk and surface periodicities are inherently mismatched. Here we reveal a hidden moiré topology emerging on the (103) surface of the Weyl semimetal NdAlSi. Angle-resolved photoemission spectroscopy uncovers closed Fermi-arc loops and momentum-space moiré modulations, phenomena unanticipated in conventional topological theory. We show that these emerge from incomplete bulk projection and multi-cell interference governed by a least-common-multiple framework. Least-common-multiple guided DFT and Green’s-function calculations quantitatively reproduce the observed spectra, establishing the universality of this commensuration rule. These findings transform a long-standing paradox of bulk-boundary correspondence into a new paradigm of momentum-space moiré reconstruction, bridging crystalline and quasicrystalline topologies and opening routes to flat-band engineering on complex surfaces.
Introduction
Since the discovery of topological materials, a central question in condensed-matter physics has been how bulk topology manifests at material boundaries. Nowhere is this connection more striking than in Weyl semimetals[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], where open Fermi arcs on high-symmetry surfaces such as (001)/(100) of tetragonal or orthorhombic lattices[5, 6, 7, 12, 13, 14] and the (0001) facet of hexagonal lattices[15, 16] directly connect the surface projections of bulk Weyl points. These canonical surfaces, where bulk and surface periodicities are trivially commensurate, have accordingly dominated both experimental and theoretical studies of topological semimetals.
In contrast, low-symmetry surfaces remain largely unexplored[17, 18, 19], despite offering a qualitatively new regime of topology. Their intrinsic incommensurability between bulk and surface Brillouin zones (SBZs) disrupts the conventional one-to-one mapping between bulk Weyl points and surface Fermi arcs, raising the question of how bulk-boundary correspondence (BBC) manifests when translational commensurability is lost. Whether BBC in this regime survives in a modified form or requires a new organizing principle remains unknown, as no direct experimental observation or coherent theoretical framework has yet addressed it. This gap stems from formidable challenges on both fronts: experimentally, low-symmetry facets rarely cleave into atomically uniform planes[19], and even when achieved, disorder and multiple terminations obscure intrinsic surface states in angle-resolved photoemission spectroscopy (ARPES) measurements; theoretically, conventional slab models assume commensurate periodicities and therefore fail to capture the correct surface repeat unit, making density functional theory (DFT) reproduction of experimental spectra notoriously difficult. Because Weyl semimetals themselves originate from broken symmetries, low-symmetry surfaces naturally extend this principle into the boundary, where reduced crystalline constraints are expected to generate qualitatively new boundary phenomena–ranging from anisotropic or reconstructed Fermi arcs[13, 20, 21] to nonlinear transport[26, 27, 28, 29] and higher-order boundary states[30]. A framework capable of reconciling bulk-surface incommensurability–both conceptually and computationally–is therefore essential for advancing this frontier.
Here, we overcome these barriers by realizing a well-defined (103) surface of the Weyl semimetal NdAlSi through precise pre-orientation and pre-cutting of single crystals, followed by controlled cleavage that exposes atomically uniform terraces. This preparation enables high-resolution ARPES measurements that reproducibly resolve multiple surface-state configurations. The spectra reveal an apparent mismatch between bulk-projected Weyl points and the SBZ–a direct manifestation of bulk-surface incommensurability. Guided by this observation, we show that the mismatch is not a breakdown of BBC but a natural outcome of incomplete bulk projection. By extending bulk-to-surface mapping beyond the first BZ, we establish a least common multiple (LCM) criterion that restores bulk-surface commensuration and explains the emergence of a “new smaller SBZ”, corresponding to the momentum-space moiré pattern observed experimentally. For NdAlSi (103), surface-projected DFT calculations based on this framework quantitatively reproduce the ARPES spectra and capture the characteristic interference among inequivalent surface cells. These results demonstrate that the observed moiré modulation is the direct consequence of LCM-guided reconstruction, where finite phase offsets between neighboring surface cells generate long-period interference and hybridization between surface and bulk-projected states. More broadly, this universal framework resolves the apparent BBC paradox and establishes low-symmetry surfaces as momentum-space analogues of moiré systems, linking incommensurability, surface reconstruction, and correlated boundary phenomena across a broad class of quantum materials.
Results
We begin with the Weyl semimetal NdAlSi[12, 13, 14, 25], a non-centrosymmetric member of the RAlSi (R = rare-earth) family[31, 32, 33, 34], which hosts well-characterized Weyl points and high-quality single crystals ideally suited for surface-sensitive spectroscopy. We focus on the low-symmetry (103) surface, which has never been explored yet provides a natural setting to probe how BBC manifests when bulk and surface periodicities are incompatible. The crystallographic structure and orientation of the (103) surface are shown in Fig. 1a, while Fig. 1b illustrates the bulk BZ and its projection. Unlike high-symmetry facets such as (001) or (100), the (103) projection exhibits an intrinsic mismatch between the periodicities of the SBZ (purple lines) and the projected bulk BZ (pink lines), as shown in Fig. 1c–this incommensurability forms the experimental starting point of our study.
The intrinsic incommensurability revealed by geometric projection motivates direct experimental verification on an atomically defined (103) surface. The surface orientation was confirmed by X-ray diffraction (XRD), which verifies the high-quality exposure of the (103) plane (Fig. 1d; see Methods for sample preparation). Fermi-surface maps acquired at 150 eV further highlight the crucial role of surface quality: uneven regions (Fig. 1e) show diffuse features dominated by bulk states, whereas well-defined terraces (Fig. 1f) exhibit sharp surface states that follow the periodicity of the surface Brillouin zone (SBZ, purple lines). A closer inspection, however, reveals an additional modulation with a smaller period (green lines), indicating a more subtle reconstruction. On the uneven region, the suppression of coherent surface states[14] exposes bulk-like bands whose periodicity does not coincide with that of the SBZ. These observations demonstrate that the mismatch between bulk and surface periodicities is not merely a geometric construct but a directly observable phenomenon on the (103) surface. This raises a central question: if Weyl points follow the bulk BZ periodicity while surface Fermi arcs adhere to the SBZ, does the BBC break down on low-symmetry surfaces?
To understand the nature of the electronic states on the (103) surface and answer the above question, we performed systematic ARPES measurements combined with bulk DFT calculations. The 3D bulk Fermi surface from DFT is shown in Fig. 2a, while Fig. 2b illustrates the (103) surface orientation and its perpendicular cross-section in momentum space. Calculated bulk Fermi surface slices along the [103] direction (k⟂) and perpendicular to the (103) surface (cyan plane in Fig. 2b) are presented in Figs. 2c and 2d, serving as theoretical benchmarks.
Experimentally, the photon energy dependent ARPES spectral intensity map at the Fermi level along ky (Fig. 2e) reveals strong k⟂ modulation, a hallmark of bulk-derived states. Systematic Fermi surface measurements over photon energies from 125 eV to 195 eV (Figs. 2f-i, 2n-q) further confirm this k⟂ dispersion. The corresponding DFT calculations (Figs. 2j-m, 2r-u) reproduce the experimental features with remarkable accuracy. More detailed bulk Fermi surfaces of the NdAlSi (103) surface are provided in Section 3 of the Supplementary Materials.
Beyond establishing the bulk band structure, these results provide critical insight into the apparent breakdown of the BBC. In particular, the evolution of bulk Fermi surface slices along [103] reveals a systematic in-plane translation of features as k⟂ is varied. Each slice produces a shifted replica of the bulk Fermi surface within the SBZ, reflecting the intrinsic periodicity of bulk states along the [103] direction. As successive bulk BZs are considered, these replicas accumulate into an effective superlattice periodicity that is distinct from the primitive bulk BZ but naturally commensurate with the SBZ.
From this perspective, the apparent mismatch between the periodicity of projected Weyl points and that of the surface Fermi arcs is a direct consequence of restricting the projection to the first bulk BZ. The bulk projections along [103] demonstrate experimentally that multiple BZs must be included: only then does the accumulated superlattice restore the periodicity needed to match the surface states. This insight provides the experimental foundation for resolving the BBC paradox, as developed below.
To clarify how this resolves the apparent BCC paradox, we track how the bulk-projection periodicity evolves as higher-order BZs are included. If only the first bulk BZ is considered, the projected Weyl points (Fig. 3a) clearly mismatch the theoretical SBZ (purple lines) and the experimental SBZ (green lines). Adding the second-order (Fig. 3b) and third-order (Fig. 3c) BZs and Weyl points along progressively restores the periodicity, bringing the bulk projection closer to the surface states. This restoration follows a LCM criterion: the projected displacement between successive bulk BZs closes only after an integer multiple of the in-plane reciprocal vectors. For the (103) surface of NdAlSi, the minimum closure occurs after steps (see Section 4 in the Supplementary Materials). As illustrated in Fig. 3d, the projected Weyl points recover the theoretical SBZ periodicity (purple lines) once higher-order BZs are included.
Beyond this restoration, the higher‐order projections reveal an emergent smaller repeating unit: a “new SBZ” (green lines in Fig. 3d) whose period along is one third of the theoretical SBZ. This “new SBZ” coincides with the one observed experimentally. Equivalently, this effect can be viewed as a fixed lateral shift between successive bulk projections (see Section 4 in the Supplementary Materials), which manifests as a momentum-space moiré pattern–a weak superlattice modulation with a period one third of the original SBZ (Fig. 3d). In general, whether a single bulk BZ suffices depends on surface symmetry. For high-symmetry surfaces such as (001) of NdAlSi, the first bulk BZ already coincides with the SBZ (Fig. 3f). For low-symmetry surfaces such as (103), however, successive bulk projections are laterally shifted (Fig. 3g), and only after steps does the bulk periodicity close. Exact spectral reconstruction is achieved only in the infinite-projection limit (see Section 4 in the Supplementary Materials). Thus, the observed mismatch is not a breakdown of the BBC but rather the natural consequence of incomplete bulk projection, while the emergent new SBZ reflects the moiré-like interference among multiple shifted projections. Here we display one representative surface-state configuration; additional terminations observed on the same cleaved surface are presented in Section 5 of the Supplementary Materials.
An additional subtlety arises when considering the surface Fermi arcs themselves. On high-symmetry surfaces such as (001) of NdAlSi, arcs from successive bulk BZs project in perfect registry, typically preserving the canonical picture of open arcs directly connecting projected Weyl points (Fig. 3f). In sharp contrast, on low-symmetry surfaces like (103), the incommensurate periodicities of bulk and SBZs cause consecutive arcs to project with lateral mismatches. This misalignment naturally drives hybridization between surface arcs originating from bulk zones that are displaced both in-plane and out-of-plane (Fig. 3g), so that the observed features are not simple replicas but collective superpositions. In certain cases, multiple arcs merge into closed SFALs–a boundary topology unusual on high-symmetry facets (Fig. 3d; see Section 6 of the Supplementary Materials for details).
Our framework, together with the best ARPES resolution presently achievable, provides compelling evidence for the emergence of SFALs. These loop-like states enrich the boundary phenomenology of Weyl semimetals: their closed -space trajectories provide natural channels for quantum interference and oscillatory responses, while the concentrated Berry curvature they enclose may enhance nonlinear or nonreciprocal transport. Moreover, overlapping arc segments can promote nesting instabilities, offering a route to correlated boundary orders. Future high-resolution probes such as spin-resolved ARPES or quantum oscillations will provide complementary confirmation and deeper insight into the microscopic structure of these loops.
To capture such arc reconstruction and hybridization theoretically, the surface periodicity used in modeling must be commensurate with bulk projections. This requirement is fulfilled by the LCM criterion, which provides a systematic prescription for selecting appropriate surface unit cells in DFT slab calculations to ensure commensurability between bulk projections and surface periodicities. In the case of the (103) surface of NdAlSi, the minimal commensuration requires , corresponding to a three-cell surface supercell composed of three consecutive unit cells. By contrast, a single-cell slab (or restricting to the first bulk BZ) systematically fails to reproduce the experimental periodicity, because the projection closure remains incomplete. With the correct three-cell supercell, surface-projected DFT captures the reduced periodicity and dominant features of the ARPES spectra. For a Si-terminated (103) surface (Fig. 4b), the LCM-consistent calculation reproduces the emergent “new smaller SBZ”, originating from the phase offsets of between bulk BZs of three consecutive, nonequivalent (103) surface unit cells. Their hybridization yields the reduced periodicity observed in experiment.
Despite this overall agreement, finer aspects of the Fermi-arc connectivity are not fully reproduced in the idealized calculations. In ARPES, closed surface Fermi-arc loops (SFALs) appear within the reduced SBZ, whereas the LCM-consistent DFT yields only open arcs. This discrepancy most likely reflects two limitations of the model: first, the absence of zigzag-type reconstructions or mixed local terminations, which in real samples enhance coupling between surface and bulk states; and second, the incomplete treatment of hybridization across multiple projected bulk BZs, whose lateral mismatch further promotes loop-like connectivity. We therefore regard the observed SFAL as an emergent property of low-symmetry surfaces beyond what is captured by ideal slab calculations. In addition to the Si-terminated regions, ARPES measurements also reveal two further surface-state configurations on other areas of the same cleaved (103) surface (see Fig. S10 in the Supplementary Materials). These distinct spectra share the reduced periodicity but do not match any of the terminations tested in Fig. S11 of the Supplementary Materials, suggesting a more complex origin such as zigzag-type reconstructions or mixed local terminations beyond the simple three-cell picture in Fig. S9b in the Supplementary Materials. A detailed discussion of these additional states, and their relation to the LCM criterion, is provided in Section 5 of the Supplementary Materials.
Beyond the ordered regions, disorder offers a complementary perspective on the same framework. When spectra are collected on a disordered area of the (103) surface, sharp surface states are strongly suppressed (Fig. 4c)[14], effectively acting as a natural “filter” that allows otherwise hidden surface bulk-projected states (SBPSs) to emerge. Remarkably, these SBPSs are well reproduced by the same surface-projected DFT calculations (Fig. 4b), underscoring the robustness of the LCM-guided modeling. When a surface state enters the bulk-projected region, it hybridizes with the bulk continuum and evolves into a surface bulk-resonant state (SBRS)[35, 36, 37, 38, 39, 40, 41, 42], producing hybridized features observable in ARPES. Momentum-resolved band dispersions along two representative cuts (Cut1 and Cut2) are shown in Figs. 4e and 4f. These sharp dispersive bands agree well with surface-projected DFT calculations (Figs. 4g and 4h). By contrast, along the same cuts on the disordered surface (Figs. 4i and 4j), broader and bulk-like bands appear, consistent with calculated SBPSs (Figs. 4k and 4l). Section 8 of the Supplementary Materials provide more details on the identification of surface states, bulk states, SBPSs, and SBRSs.
Discussion and conclusion
By investigating the low-symmetry (103) surface of the Weyl semimetal NdAlSi, we uncover a previously hidden regime of topological boundary reconstruction that fundamentally extends the conventional notion of BBC. While Weyl points follow the intrinsic bulk periodicity, the surface Fermi arcs adhere to a reduced periodicity defined by the SBZ. What initially appears as a paradox of the BBC is in fact a consequence of incomplete bulk projection, where successive bulk BZs generate laterally phase-shifted replicas that close only after multiple unit cells. This reconstruction restores commensuration through a LCM criterion, naturally producing the emergent “new smaller SBZ” observed in ARPES. The resulting interference pattern manifests as a momentum-space moiré modulation, in which neighboring surface unit cells contribute discrete phase offsets analogous to twisted or displaced layers in real-space moiré systems.
This momentum-space moiré analogy carries profound implications. The periodic phase interference between inequivalent surface unit cells can compress surface dispersions and generate flat-band-like electronic structures, thereby amplifying correlation effects at the boundary. In this picture, low-symmetry surfaces emerge as natural moiré analogues in momentum space, where the interplay of topology, symmetry, and interference gives rise to boundary states beyond the canonical open-arc paradigm. The observation of surface Fermi-arc loops (SFALs) further demonstrates that low-symmetry surfaces host boundary topologies inaccessible on conventional facets, pointing to a new class of topological states governed by multi-cell commensuration and hybridization.
More broadly, the LCM framework provides a unified prescription for reconciling bulk and surface periodicities across complex crystals. Its mathematical structure generalizes naturally to Dirac and nodal-line systems, higher-order topological phases, and non-centrosymmetric superconductors. Beyond crystalline materials, the same principle extends to quasicrystals and aperiodic systems, where commensuration becomes infinite and quasi-periodicity replaces translational order. From a practical standpoint, the LCM criterion offers a robust guideline for determining the minimal surface supercell required for accurate DFT slab calculations–long a bottleneck in modeling low-symmetry surfaces.
In summary, our work transforms a perceived inconsistency of the BBC into a new paradigm of boundary reconstruction, establishing a bridge between topological surface physics and moiré interference phenomena. By revealing that incommensurability itself can serve as a design principle for emergent boundary states, these findings open pathways toward flat-band formation, enhanced correlations, and topological interference effects in both crystalline and quasiperiodic systems.
Methods
Sample growth Single crystals of NdAlSi were grown from Al as flux. Nd, Al, Si elements were sealed in an alumina crucible with the molar ratio of 1 : 10 : 1. The crucible was finally sealed in a highly evacuated quartz tube. The tube was heated up to 1273 K, maintained for 12 hours and then cooled down to 973 K at a rate of 3 K per hour. Single crystals were separated from the flux by centrifuging. The Al flux attached to the single crystals were removed by dilute NaOH solution.
Sample preparation of the (103) surface To obtain the low-symmetry (103) surface of NdAlSi, we first determined the crystallographic orientation of the single crystals by single crystal XRD. Based on the identified orientation, the crystals were cut with a high-precision diamond wire saw to expose the desired (103) direction. Prior to mounting, a shallow laser pre-cut was introduced along the target orientation, which facilitated controlled cleavage of the sample inside the ARPES chamber. This procedure enabled reliable preparation of well-defined (103) surfaces suitable for spectroscopic measurements.
ARPES measurements High-resolution ARPES measurements were performed at the I05 beamline of the Diamond synchrotron and at the Bloch beamline of the MAX IV synchrotron. The total energy resolution (analyzer and beamline) was set at 1520 meV for the measurements. The angular resolution of the analyser was 0.1 degree. The beamline spot size on the sample was about 50 m50 m at the I05 beamline of the Diamond synchrotron and about 10 m12 m at the Bloch beamline of the MAX IV synchrotron. The samples were cleaved in situ and measured in ultrahigh vacuum with a base pressure better than 1.010-10 mbar at about 8 K at the I05 beamline of the Diamond synchrotron and about 18 K at the Bloch beamline of the MAX IV synchrotron.
DFT calculations The electronic structure calculations for NdAlSi were performed based on the density functional theory (DFT)[43, 44] as implemented in the VASP package[45, 46]. The exchange-correlation functional was treated within the generalized gradient approximation (GGA) in the form of Perdew–Burke–Ernzerhof (PBE)[47]. The projector augmented wave (PAW) method[48, 49] was used to describe the electron–ion interactions. Since the measurement temperature is above the Curie temperature, NdAlSi is non-magnetic; therefore, the Nd pseudopotential without 4f electrons was employed in the calculations. A plane-wave kinetic energy cutoff of 360 eV was used. Brillouin zone sampling was performed using a 6×6×6 Monkhorst–Pack k-point grid [50]. A Gaussian smearing width of 0.1 eV was applied to approximate the Fermi–Dirac distribution. Spin–orbit coupling (SOC) was included as a second-order perturbation.
The bulk Fermi surface of NdAlSi was computed using a 60×60×60 k-point grid including SOC. The projection of bulk states onto the plane for different -space slices was calculated via the Green’s function method, using a Fermi broadening of 0.01 eV to obtain the spectral function.
For the (103) surface states and Fermi arcs, we employed an iterative Green’s function method to model a semi-infinite surface and compute its surface states. Among the four possible terminations of the (103) surface, the Si-Al plane termination yielded results consistent with experimental observations.
Data Availability
The authors declare that all data supporting the findings of this study are available within the paper and its Supplementary Information files.
Acknowledgement
The authors thank Maxim Breitkreiz for his useful discussion of bulk-boundary correspondence. The work presented here was financially supported by the Swedish Research council (2019-00701 and 2019-03486) and the Knut and Alice Wallenberg foundation (2018.0104). Y.G.S. acknowledges the National Natural Science Foundation of China (Grants No. U2032204), the Informatization Plan of Chinese Academy of Sciences (CAS-WX2021SF-0102) and the National Key R&D Program of China 2024YFA1408400. Q. S. W. acknowledges the National Natural Science Foundation of China No. 12274436. We acknowledge MAX IV Laboratory for time on Beamline BLOCH under Proposal 20241395 and 20250284 as well as Diamond Light Source for time on Beamline I05 under Proposal SI39652. Research conducted at MAX IV, a Swedish national user facility, is supported by the Swedish Research council under contract 2018-07152, the Swedish Governmental Agency for Innovation Systems under contract 2018-04969, and Formas under contract 2019-02496.
Author Contributions
C.L. proposed and conceived the project. C.L. carried out the ARPES experiments with the assistance from M.H.B. and F.S.. Z.L.Y. contributed to the band structure calculations and theoretical discussions. H.X.L. and Y.G.S. contributed to NdAlSi crystal growth. C.L. and D.P. contributed to XRD measurements. C.L. developed the data analysis software, performed the data analysis, contributed to the theoretical derivation and wrote the paper. T.K., J.O. and B.T. provided the beamline support. C.L., Z.L.Y., M.H.B., D.P., J.F.Z., T.X., Q.S.W. and O.T. participate in the scientific discussions. O.T. revised the manuscript. All authors participated in and commented on the paper.
Competing Interests
The authors declare no competing interests.